By what number should 1365 be divided to get 31 as quotient and 32 as remainder?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
Dividend = 1365,
Quotient = 31
and Remainder = 32
Let us consider,
Divisor as a number y.

To Find: The value of y As per Euclid's lemma,
we have
Dividend = (divisor *quotient) + remainder
After substituting the values, we get

1365 = (y × 31) + 32
1365 - 32 = 31y
1333 = 31y
or y = 1333/31 = 43
So, 1365 should be divided by 43 to get 31 as quotient and 32 as remainder.

Thanks